# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ 3.1. Construction of the estimator
A key step for the generalization of the SCM to the case of outcomes residing in geodesic metric spaces is the construction of the weight vector $\overline{\bm{w}}$ in (3.2) by adopting a metric version of the weighted average $Y_{2:J+1,t}^{(\bm{w})}$ that corresponds to a weighted Fréchet mean or barycenter (Fréchet,, 1948; Ginestet et al.,, 2012). Theoretical justifications are provided in the next subsection.
3.2. Causal model for the geodesic synthetic control method
$Y_{j,t}(N)=g_{t}(U_{j}),\quad t=1,\ldots,T,\ j=1,\ldots,J+1,$ (3.3)
where $U_{j}\in\mathcal{M}$ is an unobservable time-invariant characteristic of unit $j$ and $g_{t}(\cdot):\mathcal{M}\to\mathcal{M}$ is a measurable function defined on the metric space $\mathcal{M}$. Figure 1 illustrates the causal framework for GSC with metric space-valued (random object) outcomes.
Figure 1. Illustration of the causal framework for the geodesic synthetic control method when $t\leq T_{0}$ and $t^{\prime}\geq T_{0}+1$. The circles symbolize random objects located in the geodesic metric space $\mathcal{M}$. The solid red arrow in the right lower subfigure represents the geodesic path from $Y_{1,t^{\prime}}^{(\mathrm{GSC})}$ to $Y_{1,t^{\prime}}$ that corresponds to the geodesic synthetic control estimator.
Define
$g_{t}^{(\bm{w})}(U_{2:J+1})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g_{t}(U_{j})),\quad\bm{w}\in\Delta^{J-1},$
$\mathbb{W}_{t}^{*}=\{\bm{w}\in\Delta^{J-1}:\,\,d(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1}))=0\},\quad t=1,\dots,T.$
We impose the following assumptions.
@@ 3.2. Causal model for the geodesic synthetic control method
    For each $\bm{w}\in\Delta^{J-1}$ and $t=1,\dots,T$, $g_{t}^{(\bm{w})}(U_{2:J+1})$ uniquely exists.
@@
    $\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}$ is not empty and $\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}\subset\bigcap_{t=T_{0}+1}^{T}\mathbb{W}_{t}^{*}$.
Assumption 3.1 (ii) ensures the existence of a non-empty subset $\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}$ within $\Delta^{J-1}$ such that for any $\bm{w}\in\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}$ the condition $d(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1}))=0$ holds for all $t=1,\dots,T$. This is the natural generalization of the perfect pretreatment fit assumption from equation (2) in Abadie et al., (2010) to the unique geodesic space setting that we consider here.
Assumption 3.1 can be verified in models of the form $g_{t}(U_{j})=\gamma_{\mu_{t},U_{j}}(\alpha_{t})$ with $\mu_{t}\in\mathcal{M}$ and $\alpha_{t}\in(0,1)$ in spaces such as networks (Example 2.1), Wasserstein space (Example 2.4), functional data (Example 2.5) as well as SPD matrices (Example 2.3) when the latter are equipped with the Frobenius metric, the power metric family, the Log-Euclidean metric or the Log-Cholesky metric. For the space of compositional data $\mathcal{S}_{+}^{d-1}$ as discussed in Example 2.2, Assumption 3.1 can be verified for the model $g_{t}(U_{j})=\mathrm{Exp}_{\mu_{t}}(A_{t}\mathrm{Log}_{U_{1}}(U_{j}))$, where $A_{t}$ is a $(d-1)\times(d-1)$ positive-definite matrix and $\mathrm{Exp}_{\mu}$ and $\mathrm{Log}_{\mu}$ are the Riemannian exponential and logarithmic maps at $\mu\in\mathcal{M}$, respectively. Details on verifying the assumption for these special cases are in Appendix A.
In the special case where the metric space $(\mathcal{M},d)$ corresponds to the Euclidean space $(\mathbb{R}^{q},\|\cdot\|)$ with the standard Euclidean norm, geodesics are straight lines and the causal model $g_{t}(U_{j})=\gamma_{\mu_{t},U_{j}}(\alpha_{t})$ simplifies to $g_{t}(U_{j})=\mu_{t}+\alpha_{t}(U_{j}-\mu_{t})$, which corresponds to a linear factor model without covariates and error terms. For the case where covariates are available, we introduce an augmented version in Section 3.3 to address potential biases in the GSC that may arise due to imperfect pretreatment fit.
@@
Suppose that Assumption 3.1 holds for model (3.3). Then $Y_{1,t}^{(\mathrm{GSC})}=Y_{1,t}(N)$ for each $t=T_{0}+1,\ldots,T$.
It is possible to extend model (3.3) to the case where the unobservable characteristics vary over time, i.e., $Y_{j,t}(N)=g_{t}(U_{j,t})$ for $t=1,\ldots,T$, $j=1,\ldots,J+1$, where $U_{j,t}\in\mathcal{M}$ is an unobservable characteristic of unit $j$ that may change over time $t$. In this scenario, result similar to Theorem 3.1 holds if we additionally assume that the set $\mathbf{W}^{*}_{t}=\{\bm{w}\in\Delta^{J-1}:d^{2}(g_{t}(U_{1,t}),g_{t}^{(\bm{w})}(U_{2:J+1,t}))=0\}$ satisfies a slightly modified version of Assumption 3.1 (ii).
Remark 3.1 (Geodesic synthetic control method with negative weights).
Instead of weight vectors situated in the simplex $\Delta^{J-1}$ in (3.2), one could also use more general weights
$\Sigma^{J-1}=\{\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\mathbb{R}^{J}:\sum_{j=2}^{J+1}w_{j}=1\}$
that may include negative elements. If such more general weight vectors satisfy a condition similar to Assumption 3.1 (ii), the results will remain unchanged, however the GSC with possibly negative weights may overfit the pre-treatment data. To improve out-of-sample performance, one possible approach is to select weights by minimizing a penalized objective function, as considered in Arkhangelsky et al., (2021).
@@
Suppose covariates $\{X_{j,t}\in(\mathcal{M}_{X},d_{X}):j=1,\dots,J+1,\,t=1,\dots,T\}$ are available, where the metric space $(\mathcal{M}_{X},d_{X})$ in general will differ from the outcome space $(\mathcal{M},d)$ and often will correspond to the Euclidean space $\mathbb{R}^{q}$ for a $q\geq 1$. One can then modify the dependence of the GSC weights on the covariates as follows,
$\tau_{t}^{(\mathrm{COV})}:=\gamma_{Y^{(\mathrm{COV})}_{1,t},Y_{1,t}},\quad t=T_{0}+1,\dots,T,$
where
$Y^{(\mathrm{COV})}_{1,t}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}(X)d^{2}(\nu,Y_{j,t}),$
$\overline{\bm{w}}(X)=(\overline{w}_{2}(X),\dots,\overline{w}_{J+1}(X))^{\prime}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d_{X}^{2}(X_{1,t},X_{2:J+1,t}^{(\bm{w})}),$
$X_{2:J+1,t}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d_{X}^{2}(\nu,X_{j,t}),\quad\bm{w}\in\Delta^{J-1}.$
Adopting the metric $d_{X}^{2}:=d_{1}^{2}+\dots+d_{k}^{2}$ for the space $\mathcal{M}_{X}:=\mathcal{M}_{1}\times\cdots\times\mathcal{M}_{k}$, one can also consider multiple covariates $X_{j,t}=(X_{j,t}^{(1)},\dots,X_{j,t}^{(k)})\in(\mathcal{M}_{1},d_{1})\times\cdots\times(\mathcal{M}_{k},d_{k})$. This would provide an extension of the approach of Abadie et al., (2010), who discuss the use of covariates for weight selection in the implementation of the SCM for Euclidean outcomes.
@@
In order to formally test for the presence of a causal effect, one may employ a placebo permutation test (Abadie,, 2021). In addition to the synthetic control predicted value $Y_{1,t}^{(\mathrm{GSC})}$, we compute $Y_{j,t}^{(\mathrm{GSC})}$ for $j=2,\ldots,J+1$ by applying the same algorithm as used to obtain $Y_{1,t}^{(\mathrm{GSC})}$, pretending the treatment unit is $j$. If there is an actual treatment effect only in unit $j=1$, then the distance $d(Y_{1,t},Y_{1,t}^{(\mathrm{GSC})})$ for the actual treatment unit should be among the most extreme in $\{d(Y_{j,t},Y_{j,t}^{(\mathrm{GSC})}):\,\,j=1,\ldots,J+1\}$. In this framework, the $p$-value for testing the null hypothesis of no causal effect by the intervention to unit $j=1$ is
@@
3.3. Augmented geodesic synthetic control method
@@
In some applications, $Y_{1,t}(N)$ may not be in the convex hull of control units $\{Y_{2,t},\dots,Y_{J+1,t}\}$ implying that $Y_{1,t}(N)\neq Y_{2:J+1,t}^{(\bm{w})}$ for any $\bm{w}\in\Delta^{J-1}$. In this case, the GSC estimator will be biased for the true causal effect $\tau_{t}$. To address this issue, we propose the augmented GSC, which is a regression-adjusted version of the GSC (Ben-Michael et al.,, 2021). The augmented GSC aims to reduce the bias of the GSC when covariates are available. If covariates are not available, one can instead use the geodesic synthetic difference-in-differences discussed in Section 6 below, which addresses the bias of the GSC by taking weighted averages not only over units but also over timepoints.
To construct the augmented GSC estimator, we introduce the notion of geodesic transport maps, with concrete examples provided in Remark 3.5 below.
Assumption 3.2.

### v2

@@ 3.1 Construction of the estimator
3.2 Causal model for the geodesic synthetic control method
To analyze the validity of the GSC estimator, we introduce a causal model describing the potential outcomes in the absence of treatment. In this framework, each unit $j$ is associated with a latent characteristic $U_{j}\in\mathcal{M}$, where $(\mathcal{M},d)$ is the outcome space, and the untreated outcome at time $t$ is generated as
$Y_{j,t}(N)=g_{t}(U_{j}),\quad t=1,\ldots,T,\ j=1,\ldots,J+1,$ (3)
with $g_{t}:\mathcal{M}\to\mathcal{M}$ a measurable map that evolves outcomes over time.
In the special case where $(\mathcal{M},d)$ is the Euclidean space $\mathbb{R}^{q}$ equipped with the standard Euclidean metric, geodesics are straight lines and the model simplifies to
$g_{t}(U_{j})=\mu_{t}+\alpha_{t}(U_{j}-\mu_{t}),$
which corresponds to the linear factor model often used in the Euclidean SCM literature (Abadie et al., 2010). Thus the causal model (3) generalizes the familiar Euclidean setting of SCM to a broad class of metric spaces. When covariates are available, we consider an augmented version in Section 3.3 to address potential biases that arise when the pretreatment fit is imperfect. Figure 1 provides an illustration of this framework.
Figure 1: Illustration of the causal framework for the geodesic synthetic control method. Circles represent random objects in the metric space $\mathcal{M}$. For $t\leq T_{0}$ the treated outcome $Y_{1,t}$ is matched by a weighted Fréchet mean of controls. For $t^{\prime}>T_{0}$ the estimator compares $Y_{1,t^{\prime}}^{(\mathrm{GSC})}$ with the observed $Y_{1,t^{\prime}}$, the geodesic connecting them representing the estimated causal effect.
We define the synthetic outcome for a weight vector $\bm{w}\in\Delta^{J-1}$ as
$g_{t}^{(\bm{w})}(U_{2:J+1})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g_{t}(U_{j})),$
that is, the weighted Fréchet mean of the control outcomes at time $t$. The goal of synthetic control is to find weights $\bm{w}$ such that this synthetic outcome reproduces the untreated outcome of the treated unit, $g_{t}(U_{1})$.
For each time point $t$, let
$\mathbb{W}_{t}^{*}=\bigl\{\bm{w}\in\Delta^{J-1}:d(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1}))=0\bigr\}$
denote the set of weight vectors that perfectly reconstruct $U_{1}$ at time $t$. We impose the following assumption, which is the direct analogue of the perfect pre-treatment fit assumption in Abadie et al. (2010).
@@ 3.2 Causal model for the geodesic synthetic control method
    For each $\bm{w}\in\Delta^{J-1}$ and $t=1,\ldots,T$, the weighted Fréchet mean $g_{t}^{(\bm{w})}(U_{2:J+1})$ exists uniquely.
@@
    There exists a nonempty set of weights $\bm{w}$ that achieve perfect fit in the pre-treatment period, i.e., $\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}\neq\O$ , and these same weights also reproduce $U_{1}$ in the post-treatment period, i.e., $\bigcap_{t=1}^{T_{0}}\mathbb{W}_{t}^{*}\subset\bigcap_{t=T_{0}+1}^{T}\mathbb{W}_{t}^{*}$.
Assumption 3.1(i) holds for the spaces discussed in Examples 15, and more generally for Hadamard spaces (Sturm, 2003), a broad class of non-positively curved spaces that possess unique geodesics. Assumption 3.1(ii) ensures that there is at least one set of weights $\bm{w}$ that perfectly reconstructs the treated unit in the pre-treatment period, and that this same set of weights also remains valid in the post-treatment period. This is the natural analogue of the perfect pretreatment fit condition in Abadie et al. (2010), now formulated for general geodesic spaces.
Assumption 3.1(ii) can be verified in a wide class of models. One example is
$g_{t}(U_{j})=\gamma_{\mu_{t},U_{j}}(\alpha_{t}),\quad\mu_{t}\in\mathcal{M},\ \alpha_{t}\in(0,1),$
where outcomes are obtained by moving along the geodesic from a reference point $\mu_{t}$ toward $U_{j}$ by a proportion $\alpha_{t}$. This model can be shown to hold in networks (Example 1), the Wasserstein space (Example 4), functional data (Example 5), and in the space of symmetric positive-definite matrices (Example 3) equipped with the Frobenius metric, the power metric family, the Log-Euclidean metric, or the Log-Cholesky metric. For compositional data on the sphere $\mathcal{S}_{+}^{d-1}$ (Example 2), Assumption 3.1 can be verified under models of the form
$g_{t}(U_{j})=\mathrm{Exp}_{\mu_{t}}\!\big(A_{t}\mathrm{Log}_{\mu_{t}}(U_{j})\big),$
where $A_{t}$ is a $(d-1)\times(d-1)$ positive-definite matrix, and $\mathrm{Exp}_{\mu}$ and $\mathrm{Log}_{\mu}$ denote the Riemannian exponential and logarithmic maps at $\mu\in\mathcal{M}$. Further details for these cases are provided in Section S.4 of the Supplementary Material.
@@
Suppose that Assumption 3.1 holds for model (3). Then $Y_{1,t}^{(\mathrm{GSC})}=Y_{1,t}(N)$ for each $t=T_{0}+1,\ldots,T$.
It is also possible to extend the model (3) to the case where the latent characteristics $U_{j}$ vary over time, i.e., $Y_{j,t}(N)=g_{t}(U_{j,t})$ with $U_{j,t}\in\mathcal{M}$. In this setting, a result similar to Theorem 3.1 holds provided that the sets $\mathbf{W}^{*}_{t}=\{\bm{w}\in\Delta^{J-1}:d(g_{t}(U_{1,t}),g_{t}^{(\bm{w})}(U_{2:J+1,t}))=0\}$ satisfy a modified version of Assumption 3.1(ii).
Remark 3.1 (Geodesic synthetic control with negative weights).
Instead of restricting to weights in the simplex $\Delta^{J-1}$, one may also allow weights in the affine simplex
$\Sigma^{J-1}=\{\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\mathbb{R}^{J}:\sum_{j=2}^{J+1}w_{j}=1\},$
which permits negative components. If such weights satisfy a condition analogous to Assumption 3.1(ii), identification results remain unchanged. However, negative weights can lead to overfitting of pre-treatment outcomes, and regularization or penalization, as in Arkhangelsky et al. (2021), may be needed to improve out-of-sample performance.
@@
Suppose covariates $\{X_{j,t}\in(\mathcal{M}_{X},d_{X}):j=1,\dots,J+1,\ t=1,\dots,T\}$ are available, where the covariate space $(\mathcal{M}_{X},d_{X})$ may differ from the outcome space $(\mathcal{M},d)$ and often corresponds to $\mathbb{R}^{q}$ with the Euclidean metric. In this case, the construction of the GSC weights can be modified to account for covariates. Specifically, we define
$\tau_{t}^{(\mathrm{COV})}:=\gamma_{Y^{(\mathrm{COV})}_{1,t},\,Y_{1,t}},\qquad t=T_{0}+1,\dots,T,$
where the synthetic control outcome is given by
$Y^{(\mathrm{COV})}_{1,t}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}(X)\,d^{2}(\nu,Y_{j,t}),$
and the covariate-dependent weights are chosen as
$\overline{\bm{w}}(X)\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d_{X}^{2}\!\big(X_{1,t},X_{2:J+1,t}^{(\bm{w})}\big),\quad X_{2:J+1,t}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d_{X}^{2}(\nu,X_{j,t}).$
If multiple covariates are available, say $X_{j,t}=(X_{j,t}^{(1)},\dots,X_{j,t}^{(k)})$, one can work in the product space $\mathcal{M}_{X}=\mathcal{M}_{1}\times\cdots\times\mathcal{M}_{k}$ with combined metric $d_{X}^{2}=d_{1}^{2}+\cdots+d_{k}^{2}$. This construction extends the covariate-adjusted SCM approach of Abadie et al. (2010) from Euclidean outcomes to general metric spaces.
@@
In order to formally test for the presence of a causal effect, one may employ a placebo permutation test (Abadie, 2021). In addition to the synthetic control outcome $Y_{1,t}^{(\mathrm{GSC})}$, we compute $Y_{j,t}^{(\mathrm{GSC})}$ for $j=2,\ldots,J+1$ by applying the same algorithm as used to obtain $Y_{1,t}^{(\mathrm{GSC})}$, pretending the treatment unit is $j$. If there is an actual treatment effect only in unit $j=1$, then the distance $d(Y_{1,t},Y_{1,t}^{(\mathrm{GSC})})$ for the actual treatment unit should be among the most extreme in $\{d(Y_{j,t},Y_{j,t}^{(\mathrm{GSC})}):\,\,j=1,\ldots,J+1\}$. In this framework, the $p$-value for testing the null hypothesis of no causal effect by the intervention to unit $j=1$ is
@@
3.3 Augmented geodesic synthetic control method
@@
In some applications, the untreated outcome $Y_{1,t}(N)$ may not lie in the convex hull of the controls $\{Y_{2,t},\dots,Y_{J+1,t}\}$, so that $Y_{1,t}(N)\neq Y_{2:J+1,t}^{(\bm{w})}$ for any $\bm{w}\in\Delta^{J-1}$. In this case, the GSC estimator can be biased for the true causal effect $\tau_{t}$. To address this issue, we propose an augmented version of GSC, which parallels the regression-adjusted synthetic control of Ben-Michael et al. (2021) and reduces bias when covariates are available.
The development of augmented GSC builds upon the notion of a geodesic transport map (Zhu and Müller, 2023), which ensures that the geodesic joining points $\alpha$ and $\beta$ can be naturally extended from any other point $\omega\in\mathcal{M}$. Formally, a geodesic transport map $\Gamma_{\alpha,\beta}:\mathcal{M}\to\mathcal{M}$ satisfies $\Gamma_{\alpha,\beta}(\alpha)=\beta$ and provides a natural way to “shift” points in $\mathcal{M}$ according to the displacement from $\alpha$ to $\beta$. In Euclidean space, this reduces to the familiar translation $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$, while in more general spaces it corresponds to parallel transport or its analogues. The existence of such maps is ensured in a wide range of spaces of practical interest.
Building on this idea, the augmented GSC corrects the bias of the standard GSC by transporting its synthetic prediction along a geodesic determined by pre-treatment regression adjustment. For Euclidean outcomes, the resulting estimator reduces exactly to the augmented synthetic control estimator of Ben-Michael et al. (2021), while for general geodesic spaces it provides a natural extension that preserves the same logic of bias correction.
For full details, including the formal definition of geodesic transport maps, concrete examples, and the detailed derivation of the augmented GSC estimator, we refer to Section S.1 of the Supplementary Material.
3.4 Extension to geodesic synthetic difference-in-differences
In addition to GSC and its augmented version, we also develop a geodesic extension of the synthetic difference-in-differences framework of Arkhangelsky et al. (2021). The proposed geodesic synthetic difference-in-differences (GSDID) estimator combines unit weights with time weights, thereby merging the advantages of GSC with the geodesic difference-in-differences of Zhou et al. (2025). This construction addresses bias that cannot be removed by unit reweighting alone, making use of both cross-sectional and temporal information. A notable property of GSDID is double robustness: even if either the unit-weighting scheme or the time-weighting scheme fails individually, their combination can still provide a valid adjustment.
The full methodology, including theoretical identification results and estimation details, is provided in Section S.2 of the Supplementary Material. To illustrate the proposed approach, we also include a real-data application that investigates the causal effect of the collapse of the Soviet Union on Russian mortality using age-at-death distributions and the Wasserstein metric. This example demonstrates how GSDID can capture complex causal effects for distributional outcomes in geodesic metric spaces.
4 Implementation and Simulations
4.1 Implementation details
Algorithm 1 outlines the implementation steps for the proposed geodesic synthetic control method. The optimization problem in Step 1 of Algorithm 1 requires customized approaches that depend on the specific metric and geometry of the underlying metric space. For networks (Example 1), symmetric positive-definite matrices (Example 3), distributions (Example 4), and functions (Example 5), the optimization reduces to a straightforward convex quadratic optimization problem that can be easily implemented (Stellato et al., 2020).
For the case of compositional data (Example 2), the optimization becomes more challenging, as it involves minimizing a bilevel objective where the weighted Fréchet mean $Y_{2:J+1,t}^{(\bm{w})}$ is itself an argmin problem. Due to the complexity of analytical differentiation when nonlinear Fréchet means are involved, we adopt a derivative-free constrained optimization strategy. Specifically, we use the Constrained Optimization BY Linear Approximations (COBYLA) algorithm (Powell, 1994), as implemented in the R package nloptr (Johnson, 2008). The weighted Fréchet mean involved in the objective function is obtained using the R package manifold (Dai et al., 2021). Further computational details, including explicit formulations of the optimization problems and solver strategies across different geodesic metric spaces, are provided in Section S.3 of the Supplementary Material.
Input: data $\{Y_{j,t}:j=1,\dots,J+1,t=1,\ldots,T\}$.
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})})$
where $Y_{2:J+1,t}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t})$ for $t=1,\ldots,T_{0}$;
2 $Y_{1,t}^{(\mathrm{GSC})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,t})\longleftarrow$ the synthetic control unit;
Algorithm 1 Geodesic Synthetic Control Method
4.2 Simulations for networks and symmetric positive-definite matrices
For the network outcomes, we use graph Laplacians residing in the space $\mathcal{L}$ equipped with the Frobenius metric (see Example 1). Networks are simulated via a weighted stochastic block model with 10 nodes divided into two equal-sized communities. Edges within and between communities are generated using Bernoulli distributions with probabilities specified by the edge probability matrix $\begin{pmatrix}0.75&0.1\\ 0.1&0.75\end{pmatrix}$. The weights of existing edges are generated as $\sin(0.1\pi t)+e^{-0.1t}((0.1j-0.5)^{2}-\sin(0.1\pi t))$. The graph Laplacian associated with each network is then considered to be the outcome.
For outcomes in the space of SPD matrices $\mathrm{Sym}_{m}^{+}$ equipped with the Log-Euclidean metric (see Example 3), we generate the underlying SPD matrices $\mu$ and $U$ from a Wishart distribution with 12 degrees of freedom and a scale matrix of $0.1\mathbf{I}_{10}$, where $\mathbf{I}_{10}$ denotes the $10\times 10$ identity matrix. The deterministic time trend and the unit-specific latent characteristic are constructed as $\mu_{t}=0.1t\mu$ and $U_{j}=e^{(0.1j-0.5)^{2}}U$, respectively. The parameter $\alpha_{t}$ governing the geodesic interpolation is set to $\alpha_{t}=\log(0.1(t+1))$.
Figure 2: Simulated counterfactual outcome for the treated unit had it not received the treatment $Y_{1,T}(N)$ (left) and the corresponding synthetic control outcome $Y_{1,T}^{(\mathrm{GSC})}$ (right) using the proposed method for (A) networks and (B) SPD matrices. For networks, nodes from the two communities are represented by red circles and blue squares, respectively. For SPD matrices, outcomes are visualized as heatmaps.
5 Real Data Examples
5.1 Impact of the 2011 Great East Japan Earthquake on Employment-Population Ratio by Industry
We analyze the impact of the 2011 Great East Japan Earthquake on the employment-population ratio by employment category, using data from the Research Institute of Economy, Trade and Industry (RIETI) database (https://www.rieti.go.jp/jp/database/R-JIP2021/index.html). These data provide annual employment figures by industry for each prefecture in Japan from 1994 to 2018. Employment is categorized into three broad groups: primary (agriculture, forestry, and fisheries), secondary (manufacturing, electricity, gas, water supply, waste management and construction) and tertiary (services, including wholesale/retail, finance, and healthcare). For each prefecture and each year, the employment shares (proportions) across these three categories are represented as a three-dimensional compositional vector. The component-wise square root transformation (see Example 2) maps these compositional outcomes onto the positive quadrant of the sphere $\mathcal{S}_{+}^{2}\subset\mathbb{R}^{3}$, which we endow with the geodesic metric. Unlike standard synthetic control approaches that typically analyze each compositional component separately, the proposed approach considers the entire compositional vector jointly. This holistic perspective reflects the intrinsic geometry of compositional data, thereby capturing interdependencies of employment across the three categories.
Figure 3: Ternary plot depicting employment compositions, where employment composition in the treated unit, Miyagi, is represented by circles, while employment compositions in the control prefectures are shown as triangles. Different years are distinguished by color. The synthetic controls for the post-treatment periods 2011, 2012 and 2013 are shown as hollow circles and are distinguished from the observed employment compositions for these years by a similar proportion of primary category employment, an increased proportion of secondary category employment and a reduced proportion of tertiary category employment.
5.2 Impact of the 1972 Abortion Legislation on Fertility Patterns in East Germany
The study of fertility rates has attracted substantial attention, with an increasing emphasis on analyzing age-specific fertility rates (ASFR) rather than relying solely on summary measures such as total fertility rate (Chen et al., 2017). Represented as fertility curves across reproductive ages, ASFR provides richer insights into fertility behaviors, allowing researchers to detect age-specific causal effects that aggregate statistics alone cannot capture. For example, a policy intervention might disproportionately influence fertility behaviors among younger women, information that is lost when relying solely on aggregate metrics.
In March 1972, East Germany introduced legislation permitting women to terminate pregnancies within the first twelve weeks, significantly liberalizing abortion access. Although the primary intention was reproductive autonomy, an unintended consequence was a substantial drop in fertility, with the total fertility rate decreasing notably from 2.1 children per woman in 1971 to 1.5 by 1975 (Büttner and Lutz, 1990). To gain deeper insights into how this policy affected fertility behavior across different reproductive ages, we applied the proposed GSC to ASFR curves. When viewed as functions over reproductive ages, these are functional data in the space $L^{2}$ (see Example 5).
The data for this analysis were obtained from the Human Fertility Database, which provides ASFR annually for ages 12 to 55 for each calendar year across different countries. We select East Germany as the treated unit, given its distinct policy change, and utilize a control group composed of 26 countries that did not experience comparable abortion liberalization during this period (Austria, Belgium, Bulgaria, Belarus, Canada, Switzerland, Czechia, Denmark, Spain, Estonia, Finland, France, Hungary, Iceland, Ireland, Italy, Japan, Lithuania, Netherlands, Portugal, Russia, Slovakia, Sweden, Ukraine, United Kingdom (England and Wales), and United States). We define 1965–1971 as pre-treatment periods and 1972–1975 as post-treatment periods, so that it is possible to capture both immediate and medium-term policy impacts. Figure 4 visualizes the ASFR curves for East Germany and the control countries from 1966 to 1975 using heat maps. A clear reduction in fertility rates is evident for East Germany during the post-treatment period (1972–1975) but other countries also experience reductions in fertility between 1965 and 1975.
Figure 4: Age-specific fertility rates (ASFR) for East Germany and the control countries from 1966 to 1977 shown as heat maps. For each of the countries, each column represents the ASFR across ages 12 to 55 for a given calendar year that is indicated at the bottom of the figure.
Figure 5: Heat maps of age-specific fertility rates (ASFR) for East Germany (left) and the synthetic control unit (right) from 1966 to 1975. Each column represents the ASFR across ages 12 to 55 (vertical axis) for each calendar year (horizontal axis).
6 Discussion and Concluding Remarks
An important modeling choice in this paper is to define the treatment effect as the geodesic between the potential outcomes $Y_{1,t}(N)$ and $Y_{1,t}(I)$. This extends the familiar Euclidean estimand $Y_{1,t}(I)-Y_{1,t}(N)$ to general spaces while preserving both magnitude and directional information. Most spaces of practical interest, including those considered in this paper, are unique geodesic spaces, so the effect is well defined. In cases where geodesics are not globally unique, one is typically interested in subsets where uniqueness holds—for example, the positive orthant of the hypersphere for compositional data. Alternative causal targets can also be considered, such as reporting the pair $(Y_{1,t}(I),Y_{1,t}^{(\mathrm{GSC})})$ or the scalar distance $d(Y_{1,t}(I),Y_{1,t}(N))$, in which case only a general metric structure is required. Uniqueness becomes essential, however, for the augmented GSC and GSDID extensions, where geodesic transport maps play a central role.
Practical implementation of the proposed methods also calls for suitable diagnostic tools. Standard checks from the Euclidean SCM can be adapted to the geodesic setting. These include examining pre-treatment fit by plotting $d(Y_{1,t},Y_{1,t}^{(\mathrm{GSC})})$ over $t\leq T_{0}$, considering transformations (Ferman and Pinto, 2021) such as demeaning from the Fréchet mean when fit is poor, and using backdating to detect anticipation effects. Robustness can be further assessed through leave-one-out analyses of the control units. Together, these diagnostics help evaluate the credibility of counterfactual predictions in applications.
There are several directions for future research. Given a specific type of non-Euclidean outcome one often has a choice of several metrics. For example for the case of distributional outcomes, the Wasserstein metric is popular and easy to handle for one-dimensional distributions but it is very tedious to apply it for multivariate distributions (Torous et al., 2024) and alternative metrics that are easier to implement for multivariate distributions such as the Fisher-Rao metric or sliced Wasserstein metric (Bonneel et al., 2015) may sometimes be preferable. There are also many metrics available for the space of symmetric positive-definite matrices and the space of networks. The specific choice of metric could have a bearing on the results of a causal analysis. Metric selection in general and especially in a causal context remains an open problem (Dubey et al., 2024).
Furthermore, it is clearly important to further develop approaches for inference and uncertainty quantification for synthetic control models and their generalizations. One possibility to create an appropriate asymptotic framework is to extend the results of Chernozhukov et al. (2021a, b). Another promising avenue for future research is to extend the geodesic synthetic control model to the case of a staggered adoption setting (Ben-Michael et al., 2022).
7 Data Availability Statement
The data that support the findings of this study are openly available in the Research Institute of Economy, Trade and Industry (RIETI) at https://www.rieti.go.jp/jp/database/R-JIP2021/index.html, the Human Fertility Database at www.humanfertility.org, and the Human Mortality Database at www.mortality.org.
SUPPLEMENTARY MATERIAL
S.1 Augmented Geodesic Synthetic Control Method
This section provides the formal development of the augmented geodesic synthetic control (GSC) method. While the main text offered the high-level motivation, here we present the assumptions explicitly, introduce the necessary geometric tools, and derive the estimator step by step.
The central idea behind the augmented GSC is to correct the potential bias of the standard GSC estimator by making use of geodesic transport maps (Zhu and Müller, 2023). These maps extend the displacement between two points in a metric space to any other point, thereby allowing synthetic predictions to be shifted consistently in accordance with regression adjustments. We formalize this notion in the following assumption.
Assumption S1.